# On the arithmetic and geometry of spaces $L_{m+1,n}$

### v1

@@ 3.1. Known results on spaces $L_{\lambda p,2}$
The polynomial $(Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2})$ appearing in the denominators of $\omega_{1},\omega_{2}$ is called the Moore determinant and denoted by $\Delta_{2}(Q_{1},Q_{2})$. This is the determinant of the Moore matrix $\begin{pmatrix}Q_{1}&Q_{2}\\ Q_{1}^{p}&Q_{2}^{p}\end{pmatrix}$. Its appearance is far from a coincidence: as we will see in Section 4, Moore determinants of higher order have a fundamental role in the generalizations of Pagot’s result for spaces $L_{\lambda p^{n-1},n}$. Moreover, results on Moore determinants will be helpful to simplify some proofs, even in the case of dimension 2. For this reason, we have collected the results we need on Moore determinants in the Appendix A.
@@
Let $Q_{1},Q_{2}\in k[X]$ be polynomials of degree $\lambda$ such that $\deg(iQ_{1}+jQ_{2})=\lambda$, for every $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$. Define the associated differential forms
@@
By Proposition 3.1, a pair $(Q_{1},Q_{2})$ of polynomials of degree $\lambda$ is a prompt for a space $L_{\lambda p,2}$ if, and only if, $Q_{1}$ and $Q_{2}$ have leading coefficients that are ${\mathbb{F}}_{p}$-independent, and are such that $\left((Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}\right)^{(p-1)}$ is a non-zero constant $d\in k^{\times}$. The number $c$ then needs to satisfy $d(c^{p})^{p-1}=-1$, and hence it is uniquely determined up to multiplication by a $p-1$-th root of unity.
Convention 3.5.
If we have polynomials $Q_{1}$ and $Q_{2}$ satisfying the conditions $(i)$ and $(ii)$ of Proposition 3.1, we can write $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}(X)=\sum r_{i}X^{i}$, and apply Proposition 2.14 to deduce that condition $(iii)$ is equivalent to the equalities $r_{p-1}=1$ and $r_{kp-1}=0$ for $k=2,\dots,\lambda(p-1)$. The $r_{i}$’s are polynomials in the coefficients of $Q_{1}$ and $Q_{2}$, and the equalities above will be used in Section 6 to classify certain spaces $L_{\lambda p,2}$. In order to simplify a frequently used notation, we set $R_{k}:=r_{kp-1}$ for $k=1,\dots,\lambda(p-1)$.
@@
Let $Q_{1},Q_{2}\in k[X]$ be polynomials of degree $\lambda$ with
@@
We have that $Q_{1}$ and $Q_{2}$ satisfy condition $(iii)$ of Proposition 3.1 and hence the polynomials $R_{k}$ of Convention 3.5 vanish for $k>1$. In particular, this is true for $R_{\lambda(p-1)}$, the coefficient of degree $\lambda(p-1)p-1$ of the polynomial $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}(X)$. To compute $R_{\lambda(p-1)}$, let us write $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}=X^{\lambda(p-1)p}\left[\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}\right]^{p-1}$, introduce the variable $Z=\frac{1}{X}$ and compute the coefficient of $Z$ in the expression in brackets above. We have:
$\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}$ $\equiv a^{p-1}b(1-s_{1}Z)^{p-1}(1-t_{1}Z)-b^{p}(1-t_{1}^{p}Z^{p})\mod\;Z^{2}$
$\equiv b\left(a^{p-1}(1-(t_{1}-s_{1})Z-b^{p-1}\right)\mod\;Z^{2}.$
@@
$\scriptstyle{\left(\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}\right)^{p-1}}$ $\equiv b^{p-1}\left(a^{p-1}(1-(t_{1}-s_{1})Z)-b^{p-1}\right)^{p-1}\mod Z^{2}$
$\equiv b^{p-1}\left((a^{p-1}-b^{p-1})^{p-1}+a^{p-1}(a^{p-1}-b^{p-1})^{p-2}(t_{1}-s_{1})Z\right)\mod Z^{2}.$
@@ 3.2. A new generic obstruction to the existence of spaces $L_{\lambda p,2}$
It is a result of Pagot (cf. [8, Theorème 1 and Theorème 2] ) that spaces $L_{p,2}$ and $L_{3p,2}$ exist only for $p=2$ and spaces $L_{2p,2}$ exist only for $p=2,3$. In this section, we show that there are no spaces $L_{\lambda p,2}$ if $p$ is large enough with respect to $\lambda$, vastly improving on the previously known situation. The genericity in the title of the section refers then to the fact that our result holds for all but finitely many primes once $\lambda$ is fixed. This allows for the finite remaining cases to be checked with a computer, since it suffices to check the existence of solutions of a polynomial system (see Convention 3.5). For $p=3$ and $\lambda=4,5$, this is done in Section 6.
Let us now state our result. To simplify the demonstration, we exclude from the statement the case $\lambda=1$, which has a known short proof (see [8, Théorème 2, Part 1]). By contrast, the proof in the other known cases $\lambda=2,3$ consists of several pages and the argument below consistently simplifies it.
@@
To prove the theorem, we need to recall some notation and establish two fundamental lemmas. If we have a space $L_{\lambda p,2}$ generated by a basis$(\omega_{1},\omega_{2})$ and we consider the polynomials $Q_{1}$ and $Q_{2}$ giving rise to $(\omega_{1},\omega_{2})$, we recall from the proof of Proposition 3.1 that, for every $[i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$, $P_{[i:j]}$ denotes the monic polynomial whose zeroes are those of $iQ_{2}-jQ_{1}$, and that $a$ denotes the quotient of the leading terms of $Q_{1}$ and $Q_{2}$, which satisfies $a\notin{\mathbb{F}}_{p}$. We are now ready to establish our lemmas:
@@
    Let $p>3\lambda$. Then the element $-a^{p}\in k$ is a zero of order $\geq\lambda+3$ of $R(X)$.
@@
$\begin{pmatrix}a_{\lambda}&a_{\lambda-1}&a_{\lambda-2}&\cdots&a_{1}&a_{0}&0&\cdots&0\\ 0&a_{\lambda}&a_{\lambda-1}&\cdots&a_{2}&a_{1}&a_{0}&\cdots&0\\ \vdots&\vdots&\vdots&&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&0&a_{\lambda}&a_{\lambda-1}&\cdots&a_{0}\\ \lambda a_{\lambda}&(\lambda-1)a_{\lambda-1}&(\lambda-2)a_{\lambda-2}&\cdots&a_{1}&0&0&\cdots&0\\ 0&\lambda a_{\lambda}&(\lambda-1)a_{\lambda-1}&\cdots&2a_{2}&a_{1}&0&\cdots&0\\ \vdots&\vdots&\vdots&&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\lambda a_{\lambda}&(\lambda-1)a_{\lambda-1}&(\lambda-2)a_{\lambda-2}&\cdots&a_{1}\end{pmatrix}.$
Since both $a_{\lambda}$ and $a_{\lambda-1}$ are independent of $t$ (the former being 1 and the latter as a result of Lemma 3.6), the first two columns of this matrix are always the same for every $t$. As a result, we are left with at most $32\lambda-$ 3 rows that contain quantities of the form $\frac{p_{ij}(t)}{a+t}$, where $p_{ij}(t)$ is either zero or a polynomial of degree 1. Applying Leibniz formula for the determinant, we then obtain that $\mathrm{Disc}(P_{t})=\frac{R(t)}{(a+t)^{2\lambda-3}}$ for some $R(t)$ of degree at most $32\lambda-$ 3 as desired.
Proof of (ii): Consider the differential form $\omega_{1}=\frac{Q_{2}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$. Remark that the poles of $\omega_{1}$ are the zeroes of the polynomial $Q_{1}^{p}-Q_{2}^{p-1}Q_{1}=\prod_{j=0}^{p-1}\left(Q_{1}-jQ_{2}\right)$. Hence, for every pole $x$ of $\omega_{1}$, there exists a number $i\in\{0,\dots,p-1\}$ such that $x$ is a zero of $iQ_{2}-Q_{1}$. The residue of $\omega_{1}$ at $x$ can then be computed as follows:
$\mathrm{res}_{\omega_{1}}(x)=\frac{-1}{\left(Q_{1}^{p}-Q_{2}^{p-1}Q_{1}\right)^{{}^{\prime}}(x)}=\frac{1}{(iQ_{2}-Q_{1})^{\prime}(x)\prod_{j\neq i}(jQ_{2}-Q_{1})(x)}=\frac{1}{(iQ_{2}-Q_{1})^{\prime}(x)(-Q_{2}^{p-1}(x))}.$
For every $i=0,\dots,p-1$, we recall from the proof of Proposition 3.1 that $iQ_{2}-Q_{1}=c(a+i)P_{[i:1]}$, and we observe that $P_{[i:1]}$ coincides with $P_{i}$ as in the statement of the Lemma. We can then rewrite the identity above as
$\mathrm{res}_{\omega_{1}}(x)=-\frac{1}{c(a+i){P_{i}^{\prime}}(x)Q_{2}(x)^{p-1}}.$
We can then consider the product $H_{i}$ of the residues at all the poles that are roots of $P_{i}$
$H_{i}:=\prod_{x\in Z(P_{i})}\mathrm{res}_{\omega_{1}}(x)=\frac{(-1)^{\lambda}}{c^{\lambda}(a+i)^{\lambda}\prod_{x\in Z(P_{i})}P_{i}^{{}^{\prime}}(x)\prod_{x\in Z(P_{i})}Q_{2}^{p-1}(x)}.$
@@
$H_{i}=\frac{(c(a+i))^{\lambda(p-2)}}{(-1)^{\frac{\lambda(\lambda+1)}{2}}\mathrm{Disc}(P_{i})\mathrm{Res}(iQ_{2}-Q_{1},Q_{2})^{p-1}}=(-1)^{\frac{\lambda(\lambda+1)}{2}}\frac{(c(a+i))^{\lambda(p-2)}}{\mathrm{Disc}(P_{i})\mathrm{Res}(-Q_{1},Q_{2})^{p-1}}.$
By Lemma 3.8 (i), we can express the discriminant $\mathrm{Disc}(P_{i})$ in terms of the polynomial $R(X)$ to obtain that $H_{i}=\delta\frac{(a+i)^{\lambda p-3}}{R(i)}$, where we have set $\delta:=(-1)^{\frac{\lambda(\lambda+1)}{2}}\frac{c^{\lambda(p-2)}}{\mathrm{Res}(-Q_{1},Q_{2})^{p-1}}$ for ease of notation, since this is independent of $i$.
Since the differential form $\omega_{1}$ is logarithmic, we have that $H_{i}\in{\mathbb{F}}_{p}^{\times}$ and in particular that $H_{i}^{p-1}=1$. The following equations then hold for every $i\in\{0,\dots,p-1\}$:
$\delta^{p-1}(a+i)^{(\lambda p-3)(p-1)}$ $=R(i)^{p-1}$
$\delta^{p-1}(a+i)^{(\lambda p^{2}-(\lambda+3)p+3)}R(i)$ $=R(i)^{p}$
$\delta^{p-1}(a+i)^{\lambda p^{2}}(a+i)^{3}R(i)$ $=R(i)^{p}(a+i)^{(\lambda+3)p}$
$\delta^{p-1}(a^{p^{2}}+i)^{\lambda}(a+i)^{3}R(i)$ $=R_{p}(i)(a^{p}+i)^{(\lambda+3)},$
@@
$\delta^{p-1}(a^{p^{2}}+i)^{\lambda}(a+i)^{3}R(i)-R_{p}(i)(a^{p}+i)^{(\lambda+3)}=0,$
which is a polynomial equation of degree at most $3\lambda$ in $i$ that is satisfied for every $i=0,\dots,p-1$. Since we have that $p>3\lambda$, this is actually an equality of univariate polynomials, namely we have
@@
in the ring $k[X]$. The right hand side of the equation admits $-a^{p}$ as root of order at least $\lambda+3$. Since $a^{p}\neq a$ and $a^{p}\neq a^{p^{2}}$ we conclude that $R(X)$ has $-a^{p}$ as root of order at least $\lambda+3$, as well. ∎
@@
In spite of its fairly elementary proof, Lemma 3.8 (i) is already quite powerful. Combining the lower bound on the order of $-a^{p}$ as a zero of $R(X)$ given in $(ii)$ and the upper bound on the degree of $R(X)$ given by $(i)$ one gets that $32\lambda-3\geq\lambda+$ 3, which gives $\lambda\geq 6$. It follows that no further argument is needed to prove Theorem 3.7 when $\lambda\leq 5$.
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $P,Q$ be coprime polynomials in $k[X]$ such that $0\leq\deg(Q)<\deg(P)<p$ and consider the polynomial function
$z\mapsto D(z):=\mathrm{Disc}(P-zQ)\in k[z].$
Let $Z_{D}$ be the set of zeroes of $D$ and $Z_{\delta}$ be the set of zeroes of $\left(\frac{P}{Q}\right)^{\prime}$ where in both sets zeroes are counted with multiplicity. Then the correspondence
$F:Z_{\delta}$ $\to Z_{D}$
$x$ $\mapsto\frac{P(x)}{Q(x)}$
is a well defined bijective function.

### v2

@@ 3.1. Known results on spaces $L_{\lambda p,2}$
It follows from the proposition that, given a pair $Q_{1},Q_{2}$ satisfying conditions $(i),(ii)$ and $(iii)$, then $Q_{1}$ and $Q_{2}$ are necessarily coprime, and have simple roots. In fact, the poles of $\omega_{1}$ being simple, we have that $\prod_{i=0}^{p-1}(Q_{1}+iQ_{2})=Q_{1}^{p}-Q_{1}Q_{2}^{p-1}$ has simple roots. In particular, $Q_{1}$ has simple roots and no roots in common with $Q_{1}+Q_{2}$, hence no roots in common with $Q_{2}$. Similarly, from the fact that $\omega_{2}$ has simple poles one derives that $Q_{2}$ has simple roots.
Remark 3.4.
The polynomial $(Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2})$ appearing in the denominators of $\omega_{1},\omega_{2}$ is called the Moore determinant and denoted by $\Delta_{2}(Q_{1},Q_{2})$. This is the determinant of the Moore matrix $\begin{pmatrix}Q_{1}&Q_{2}\\ Q_{1}^{p}&Q_{2}^{p}\end{pmatrix}$. Furthermore, the polynomial $\frac{Q_{2}^{p^{2}-1}-Q_{1}^{p^{2}-1}}{Q_{2}^{p-1}-Q_{1}^{p-1}}$ is the Dickson invariant $c_{2,1}$ evaluated in $Q_{1},Q_{2}$. The appearance of Moore determinants and Dickson invariants is far from a coincidence: as we will see in Section 4, these have a fundamental role in the generalizations of Proposition 3.1 for spaces $L_{\lambda p^{n-1},n}$. Moreover, Moore determinants will be helpful to simplify some proofs, even in the case of dimension 2. For this reason, we have collected the results we need on Moore determinants and Dickson invariants in Appendix A.
@@
Let $Q_{1},Q_{2}\in k[X]$ be polynomials of degree $\lambda$ such that $\deg(iQ_{1}+jQ_{2})=\lambda$, for every $i:j]\in{\mathbb{P}}^{1}({\mathbb{F}}_{p})$. Define the associated differential forms
@@
By Proposition 3.1, a pair $(Q_{1},Q_{2})$ of polynomials of degree $\lambda$ is a prompt for a space $L_{\lambda p,2}$ if, and only if, $Q_{1}$ and $Q_{2}$ have leading coefficients that are ${\mathbb{F}}_{p}$-independent, and are such that $\left((Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}\right)^{(p-1)}$ is a non-zero constant $d\in k^{\times}$. The number $c$ then needs to satisfy $d(c^{p})^{p-1}=-1$, and hence it is uniquely determined up to multiplication by a p1p-1-th root of unity.
Convention 3.7.
If we have polynomials $Q_{1}$ and $Q_{2}$ satisfying the conditions $(i)$ and $(ii)$ of Proposition 3.1, we can write $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}(X)=\sum r_{i}X^{i}$, and apply Proposition 2.15 to deduce that condition $(iii)$ is equivalent to the equalities $r_{p-1}=1$ and $r_{kp-1}=0$ for $k=2,\dots,\lambda(p-1)$. The $r_{i}$’s are polynomials in the coefficients of $Q_{1}$ and $Q_{2}$, and the equalities above will be used in Sections 6 and 7 to classify certain spaces $L_{\lambda p,2}$. In order to simplify a frequently used notation, we set $R_{k}:=r_{kp-1}$ for $k=1,\dots,\lambda(p-1)$.
@@
Assume $(\lambda,p)\neq(1,2)$ and let $Q_{1},Q_{2}\in k[X]$ be polynomials of degree $\lambda$ with
@@
We have that $Q_{1}$ and $Q_{2}$ satisfy condition $(iii)$ of Proposition 3.1 and hence the polynomials $R_{k}$ of Convention 3.7 vanish for $k>1$. In particular, this is true for $R_{\lambda(p-1)}$, the coefficient of degree $\lambda(p-1)p-1$ of the polynomial $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}(X)$. To compute $R_{\lambda(p-1)}$, let us write $(Q_{1}^{p}-Q_{1}Q_{2}^{p-1})^{p-1}=X^{\lambda(p-1)p}\left[\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}\right]^{p-1}$, introduce the variable $Z=\frac{1}{X}$ and compute the coefficient of $Z$ in the expression in brackets above. We have:
$\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}$ $\equiv a^{p-1}b(1-s_{1}Z)^{p-1}(1-t_{1}Z)-b^{p}(1-t_{1}^{p}Z^{p})\mod\;Z^{2}k[Z]$
$\equiv b\left(a^{p-1}(1-(t_{1}-s_{1})Z-b^{p-1}\right)\mod\;Z^{2}k[Z].$
@@
$\scriptstyle{\left(\left(\frac{Q_{1}}{X^{\lambda}}\right)^{p-1}\frac{Q_{2}}{X^{\lambda}}-\left(\frac{Q_{2}}{X^{\lambda}}\right)^{p}\right)^{p-1}}$ $\equiv b^{p-1}\left(a^{p-1}(1-(t_{1}-s_{1})Z)-b^{p-1}\right)^{p-1}\mod Z^{2}k[Z]$
$\equiv b^{p-1}\left((a^{p-1}-b^{p-1})^{p-1}+a^{p-1}(a^{p-1}-b^{p-1})^{p-2}(t_{1}-s_{1})Z\right)\mod Z^{2}k[Z].$
@@ 3.2. A new generic obstruction to the existence of spaces $L_{\lambda p,2}$
It is a result of Pagot (cf. [Pagot02, Theorème 1 and Theorème 2]) that spaces $L_{p,2}$ and $L_{3p,2}$ exist only for $p=2$ and spaces $L_{2p,2}$ exist only for $p=2,3$. In this section, we show that there are no spaces $L_{\lambda p,2}$ if $p$ is large enough with respect to $\lambda$, vastly improving on the previously known situation. The genericity in the title of the section refers then to the fact that our result holds for all but finitely many primes once $\lambda$ is fixed, which allows for a direct computation of the remaining cases. For example, for $\lambda=4$ we only need to check the primes $p\in\{2,3,5,7,11\}$. For $p=2$ the result is classical (see Remark 4.29), the case $p=3$ is achieved in Section 6.1 and the case $p\in\{5,7,11\}$ is achieved in Section 7, completing the classification of spaces $L_{4p,2}$.
Let us now state our result. To simplify the demonstration, we exclude from the statement the case $\lambda=1$, which has a known short proof (see [Pagot02, Théorème 2, Part 1]). By contrast, the proof in the other known cases $\lambda=2,3$ consists of several pages and the argument below consistently simplifies it.
@@
[…]
@@
    Let $p>3\lambda$. Then the element $a^{p}\in k$ is a root of $R(X)$ of order $\geq\lambda+3$.
@@
$\begin{pmatrix}a_{\lambda}&a_{\lambda-1}&\cdots&a_{2}&a_{1}&a_{0}&\cdots&0\\ 0&a_{\lambda}&\cdots&a_{3}&a_{2}&a_{1}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&&\vdots\\ 0&0&\cdots&a_{\lambda}&a_{\lambda-1}&a_{\lambda-2}&\cdots&a_{0}\\ \lambda a_{\lambda}&(\lambda-1)a_{\lambda-1}&\cdots&2a_{2}&a_{1}&0&\cdots&0\\ 0&\lambda a_{\lambda}&\cdots&3a_{3}&2a_{2}&a_{1}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\lambda a_{\lambda}&(\lambda-1)a_{\lambda-1}&(\lambda-2)a_{\lambda-2}&\cdots&a_{1}\end{pmatrix}$.
Since both $a_{\lambda}$ and $a_{\lambda-1}$ are independent of $t$ (the former being 1 and the latter as a result of Lemma 3.8), the first two columns of this matrix are always the same for every $t$. As a result, we are left with at most $2\lambda-3$ rows that contain quantities of the form $\frac{p_{ij}(t)}{a+t}$, where $p_{ij}(t)$ is either zero or a polynomial of degree 1. Applying Leibniz formula for the determinant, we then obtain that $\mathrm{Disc}(P_{t})=\frac{R(t)}{(a+t)^{2\lambda-3}}$ for some $R(t)$ of degree at most $2\lambda-3$ as desired.
Proof of (ii): We consider the differential form $\omega_{2}=\frac{-Q_{1}dX}{Q_{1}Q_{2}^{p}-Q_{1}^{p}Q_{2}}$ and remark that the poles of $\omega_{2}$ are the zeroes of the polynomial $Q_{2}^{p}-Q_{1}^{p-1}Q_{2}=\prod_{j=0}^{p-1}\left(Q_{2}-jQ_{1}\right)$. Hence, for every pole $x$ of $\omega_{2}$, there exists a number $j\in\{0,\dots,p-1\}$ such that $x$ is a zero of $Q_{2}-jQ_{1}$. The residue of $\omega_{2}$ at $x$ can then be computed as follows:
$\mathrm{res}_{\omega_{2}}(x)=\frac{-1}{\left(Q_{2}^{p}-Q_{1}^{p-1}Q_{2}\right)^{{}^{\prime}}(x)}=\frac{-1}{(Q_{2}-jQ_{1})^{\prime}(x)\prod_{i\neq j}(Q_{2}-iQ_{1})(x)}=\frac{1}{(Q_{2}-jQ_{1})^{\prime}(x)(Q_{1}^{p-1}(x))}.$
For every $j=0,\dots,p-1$, we recall from the proof of Proposition 3.1 that $Q_{2}-jQ_{1}=c(a+j)P_{[1:j]}$, and we observe that $P_{[1:j]}$ coincides with the polynomial $P_{j}$ defined in the statement of the Lemma. We can then rewrite the identity above as
$\mathrm{res}_{\omega_{2}}(x)=\frac{1}{c(a+j){P_{j}^{\prime}}(x)Q_{1}(x)^{p-1}}.$
Considering the product $H_{j}$ of the residues at all the poles that are roots of $P_{j}$, we have that
$H_{j}:=\prod_{x\in Z(P_{j})}\mathrm{res}_{\omega_{2}}(x)=\frac{1}{c^{\lambda}(a+j)^{\lambda}\prod_{x\in Z(P_{j})}P_{j}^{{}^{\prime}}(x)\prod_{x\in Z(P_{j})}Q_{1}^{p-1}(x)}.$
@@
$H_{j}=\frac{1}{(-1)^{\frac{\lambda(\lambda-1)}{2}}\mathrm{Disc}(P_{j})(c(a+j))^{-\lambda(p-2)}\mathrm{Res}(Q_{2}-jQ_{1},Q_{1})^{p-1}}=(-1)^{\frac{\lambda(\lambda-1)}{2}}\frac{(c(a+j))^{\lambda(p-2)}}{\mathrm{Disc}(P_{j})\mathrm{Res}(Q_{2},Q_{1})^{p-1}}.$
By Lemma 3.10 (i), we can express the discriminant $\mathrm{Disc}(P_{j})$ in terms of the polynomial $R(X)$ to obtain that $H_{j}=\delta\frac{(a+j)^{\lambda p-3}}{R(j)}$, where we have set $\delta:=(-1)^{\frac{\lambda(\lambda-1)}{2}}\frac{c^{\lambda(p-2)}}{\mathrm{Res}(Q_{2},Q_{1})^{p-1}}$ for ease of notation, since this is independent of $j$.
Since the differential form $\omega_{2}$ is logarithmic, we have that $H_{j}\in{\mathbb{F}}_{p}^{\times}$ and in particular that $H_{j}^{p-1}=1$. The following equations then hold for every $j\in\{0,\dots,p-1\}$:
$\delta^{p-1}(a+j)^{(\lambda p-3)(p-1)}$ $=R(j)^{p-1}$
$\delta^{p-1}(a+j)^{(\lambda p^{2}-(\lambda+3)p+3)}R(j)$ $=R(j)^{p}$
$\delta^{p-1}(a+j)^{\lambda p^{2}}(a+j)^{3}R(j)$ $=R(j)^{p}(a+j)^{(\lambda+3)p}$
$\delta^{p-1}(a^{p^{2}}+j)^{\lambda}(a+j)^{3}R(j)$ $=R_{p}(j)(a^{p}+j)^{(\lambda+3)},$
@@
$\delta^{p-1}(a^{p^{2}}+j)^{\lambda}(a+j)^{3}R(j)-R_{p}(j)(a^{p}+j)^{(\lambda+3)}=0,$
which is a polynomial equation of degree at most $3\lambda$ in $j$ that is satisfied for every $j=0,\dots,p-1$. Since we have that $p>3\lambda$, this is actually an equality of univariate polynomials, namely we have that
@@
holds in the ring $k[X]$. The right hand side of the equation admits $a^{p}$ as root of order at least $\lambda+3$. Since $a^{p}\neq a$ and $a^{p}\neq a^{p^{2}}$ we conclude that $R(X)$ has $a^{p}$ as root of order at least $\lambda+3$, as well. ∎
@@
Despite its fairly elementary proof, Lemma 3.10 (i) is quite powerful. Combining the lower bound on the order of $a^{p}$ as a zero of $R(X)$ given in $(ii)$ and the upper bound on the degree of $R(X)$ given by $(i)$ one gets that $2\lambda-3\geq\lambda+3$, which gives $\lambda\geq 6$. It follows that no further argument is needed to prove Theorem 3.9 when $\lambda\leq 5$. In order to prove it in full generality, we need to go a bit further.
In order to prove the auxiliary lemma below, we introduce the following notation: for polynomials $A,B\in k[X]$, we denote by $A\wedge B\in k[X]$ the monic gcd of $A$ and $B$ and by $\mathrm{rdeg}(A)$ the degree of the reduction of $A$, that is, the number of distinct roots of the polynomial $A$. Finally, if $A=\prod_{i=1}^{\mathrm{rdeg}(A)}(X-x_{i})^{m_{i}}$, we denote by $Z_{A}=\{x_{1},\dots,x_{\mathrm{rdeg}{A}}\}$ the set of zeroes of $A$, and we define $\hat{\prod}_{x\in Z_{A}}B(x):=\prod_{i=1}^{\mathrm{rdeg}{A}}B(x_{i})^{m_{i}}$.
@@
Let $k$ be an algebraically closed field of characteristic $p>0$. Let $P,Q$ be coprime polynomials in $k[X]$ such that $0\leq\deg(Q)<\deg(P)<p$ and let $D(Z)\in k[Z]$ be the discriminant in degree $\deg(P)$ of the polynomial $P(X)-ZQ(X)\in k[X]$.
Then, the derivative $\left(\frac{P}{Q}\right)^{\prime}$ of the rational function $\frac{P}{Q}\in k(X)$ can be written as
$\left(\frac{P(X)}{Q(X)}\right)^{\prime}=\frac{A(X)}{B(X)Q(X)}$
such that $A\wedge Q=1$ and $A(X)=\prod_{i=1}^{d}(X-x_{i})$ with $x_{1},\dots,x_{d}\in k$ and $d=\deg(P)+\mathrm{rdeg}(Q)-1$. Moreover, the $x_{i}$’s are such that
$D(Z)=c\prod_{i=1}^{d}\left(Z-\frac{P(x_{i})}{Q(x_{i})}\right)\;\mbox{for some}\;c\in k^{\times}.$
In particular, $\deg(D)=\deg(P)+\mathrm{rdeg}(Q)-1$.